OpenStudy (terminator9999):

Use the law of cosines to find the value of 2*3*4 cos. WILL GIVE MEDALS!!!! (Link to pic below)

OpenStudy (terminator9999):

The different answers are -24 21 -21 24

OpenStudy (mysticfury):

would this help? http://www.mathwarehouse.com/trigonometry/law-of-cosines-formula-examples.php

OpenStudy (wolf1728):

Go here: http://www.1728.org/trigtut2.htm and you'll find this formula:

OpenStudy (wolf1728):

I figured you'd need another graphic too. cos (A) = (b^2 + c^2 -a^2) / 2 * b *c cos (A) = (4^2 + 3^2 -2^2) / 2 * 4 * 3 cos (A) = (16 + 9 -4) / 24 cos (A) = 21 / 24 cos (A) = .875 arc Cos (.875) = 28.955 Degrees

OpenStudy (wolf1728):

Your question asks "find the value of 2*3*4 cos". I'm guessing cos means cos theta = cos(28.955) which equals .875 So, 2*3*4 * .875 = 21

OpenStudy (anonymous):



OpenStudy (wolf1728):

terminator9999 Did you ever figure out an answer to this?